OpenStudy (anonymous):

Compare: two-ninths to one-fourth less than greater than equal to None of the above All of the above

OpenStudy (anonymous):

\[\frac{2}{9}=\frac{2}{9}\times \frac{4}{4}=\frac{8}{36}\] \[\frac{1}{4}=\frac{1}{4}\times \frac{9}{9}=\frac{9}{36}\] now it should be easy to compare them

OpenStudy (anonymous):

so its = 2

OpenStudy (anonymous):

@No_One_Knows_me